OpenStudy (anonymous):

3/8 times 40 = ? 3 15 5 25

OpenStudy (anonymous):

3/8*40/1=(3*40)/(8*1)=120/8=15

OpenStudy (anonymous):

\[\huge \color{green}{\frac{3}{8} \times 40 = \frac{3}{\cancel{8}} \times \cancel{40}^5 = 3 \times 5 = 15}\]